2) Following through on your understanding of a logarithm being an exponent,
what would log b(ba) equal?
and
what would 10 log(x) equal?

Answer by stanbon(75887)

2) Following through on your understanding of a logarithm being an exponent,
what would log b(ba) equal?
logb(ab) = logb(a) + logb(b) = logb(a) + 1
and
what would 10 log(x) equal?
10 log(x) = log(x^10)
================
Cheers,
Stan H.